# Burnside type results for fusion categories

### v10

@@ 7.3. Computations of the regular components
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. With the above notations, for all $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$,
(7.4) $F_{j}=\frac{1}{n(H)}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}\big{)}.$
@@
For any $j$ such that $\mu_{j}\in{G(\widehat{H})}$,
$F_{j}$ $=$ $\frac{1}{n_{j}}\big{(}\sum_{i=1}^{m}h_{i}\mu_{j}(x_{i^{*}})x_{i}\big{)}=\frac{1}{n_{j}}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\sum_{x_{i}\in\mathcal{B}_{\psi}}\big{(}h_{i}\mu_{j}(x_{i^{*}})x_{i}\big{)}=$
$=$ $\frac{1}{n_{j}}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi^{-1}(\mu_{j})R_{\psi}\big{)}=\frac{1}{n(H)}\big{(}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}\big{)}.$
If $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ then $n_{j}=n(H)$ by Corrolary 2.22. ∎
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. With above notations, for any linear character $\psi$ of ${G(\widehat{H})}$, the set $\mathcal{B}_{\psi}$ is not empty.
@@
Previous corollary shows that the linear span of $<F_{j}\;|\;\mu_{j}\in{G(\widehat{H})}>$ lies inside the span of $<\mathcal{B}_{\psi}\;|\;\mathcal{B}_{\psi}\neq\emptyset>$. By counting the dimension of these vector spaces, it follows that $\mathcal{B}_{\psi}$ is not empty, for any $\psi$ in the dual. ∎
From Equation (7.4), it follows immediately:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian hypergroup satisfying any of the two hypothesis from Lemma 2.23. Then
$R_{\psi}=\frac{1}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}\psi(\mu_{j})F_{j}\big{)}.$
Proof of Theorem 7.8
By Lemma 7.15, if $\mu_{j}=\mu_{1}$, then
$R_{1}=\frac{1}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}F_{j}\big{)}=\lambda_{{H_{ad}}}.$
Corollary 6.10 proves that $\mathcal{B}_{1}=\mathcal{B}_{ad}$. Therefore the above grading
@@
coincides with the universal grading of ${(H,\mathcal{B})}$. Thus $U_{H}={G(\widehat{H})}$. By duality, $U_{\widehat{H}}=G(H)$ if ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is abelian.
7.4. On the perp of the adjoint subring
Recall from Definition 1.6 that an abelian RN-hypergroup is called dualizable if ${\widehat{(H,\mathcal{B})}}$ is also RN.
For any sub-hypergroup ${(L,\;\mathcal{S})}$ of an abelian dualizable hypergroup ${(H,\mathcal{B})}$, define
@@ 7.4. On the perp of the adjoint subring
With the above notations, $\mathcal{S}^{\perp}=\bigcap_{s\in\mathcal{S}}\ker_{(H,\mathcal{B})}(s)$. Moreover, [Har79, Proposition 2.11] implies that $\mathcal{S}^{\perp}$ is a sub-hypergroup of ${\widehat{(H,\mathcal{B})}}$, and $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$.
@@
Let $H$ be an abelian dualizable RN-hypergroup. With the above notations,
@@
Consider the trivial character $\psi_{1}=\mu_{1}\big{|}_{G(\widehat{H})}$. It follows that
@@
By duality, one obtains the other equality. ∎
@@ 7.5. Quotients of hypergroups
In this subsection, we define the quotient hypergroup of a RN-hypergroup ${(H,\mathcal{B})}$ by a sub-hypergroup ${(L,\;\mathcal{S})}$. This is the analogue of the quotient construction from [Har79], in the case of probability groups. Recall that a probability group is a symmetric normalized RN-hypergroup.
One can define the following equivalence relation on $\mathcal{B}$. For two elements $a,b\in\mathcal{B}$, say $a\sim_{\mathcal{S}}b$ if and only if there are $s_{1},s_{2}\in{\mathcal{S}},x\in{\mathcal{B}}$ such that $m(x,as_{1})>0$ and $m(x,s_{2}b)>0$.
Let $[a]_{\mathcal{S}}$ denote the equivalence class of any $a\in\mathcal{B}$, with respect to $\sim_{\mathcal{S}}$. Define $H//{\mathcal{S}}$ as the set of all these equivalence classes of $\sim_{\mathcal{S}}$. Recall that $\lambda_{\mathcal{S}}$ is the primitive central idempotent of ${\mathcal{S}}$ corresponding to $\mu_{1}=\mathrm{FPdim}$.
For an element $a\in\mathcal{B}$, one can see that $[a]_{\mathcal{S}}=[b]_{\mathcal{S}}$ if and only if $\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}=\lambda_{\mathcal{S}}b\lambda_{\mathcal{S}}$. Therefore there is a set bijection
@@
Then, it easy to verify that $H//{\mathcal{S}}$ becomes a ARN-hypergroup, with the multiplication inherited from $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, via the above isomorphism. We denote by $\overline{m}({[c]},\;[a][b])$ the multiplicity structure of $H//{\mathcal{S}}$. Therefore,
@@
We write shortly $[a]$ instead of $[a]_{\mathcal{S}}$ when no confusion is possible. Following [Har79], if ${(H,\mathcal{B})}$ is abelian, one can show that
@@ 7.6. Applying Harrison’s results
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. In this subsection, we also use the short notation $H$ (or $\mathcal{B}$) for this hypergroup, if $\mathcal{B}$ (or $H$) are implicitly understood. For any sub-hypergroup ${(L,S)}$ of an ANR-hypergroup ${(H,\mathcal{B})}$, note that [Har79, Proposition 2.11] implies that there is an isomorphism of hypergroups:
@@
Dually, for ${\mathcal{S}}={(\widehat{H})_{ad}}$, one obtains that
@@ 8. On lower and upper central series of hypergroups
Let ${(L,\;\mathcal{S})}$ be a sub-hypergroup of an abelian RN-hypergroup ${(H,\mathcal{B})}$. We define the commutator ${(L,\;\mathcal{S})}^{{co}}$ of ${(L,\;\mathcal{S})}$ in ${(H,\mathcal{B})}$ as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by all standard elements $x\in\mathcal{B}$ such that $xx^{*}\in\mathcal{{\mathcal{S}}}$.
@@
for any sub-hypergroup $\mathcal{S}$ of $\mathcal{B}$. By applying $\perp$ to (8.1), one obtains
@@
For an ARN-hypergroup ${(H,\mathcal{B})}$, define $H_{(0)}=\mathbb{C}$ and $H_{(n)}=(H_{(n-1)})^{{co}}$, for all $n\geq 1$. Then, $H_{(1)}=\mathbb{C}^{{co}}=G({(H,\mathcal{B})})=H_{pt}$.
@@
Similarly to [GN08, Theorem 4.16], one can prove that $H^{(n)}=\mathbb{C}$ if and only if $H_{(n)}=H$, where $H$ is an ARN-hypergroup.
@@
An ARN-hypergroup is called nilpotent when $H^{(n)}=\mathbb{C}$, for some $n\in\mathbb{N}$.
@@
It reduces to show that $\mu_{j}\mu_{j^{\#}}\in({\mathcal{S}}^{co})^{\perp}$, for all $\mu_{j}\in{\mathcal{S}}^{\perp}$.
@@
(8.4) $\widehat{H}^{(n)}\subseteq H_{(n)}^{\perp}$
@@
$\widehat{H}^{(0)}=H=\mathbb{C}^{\perp}=H_{(0)}^{\perp}.$
@@
By duality, one obtains the second inclusion. ∎
@@
If ${(H,\mathcal{B})}$ is nilpotent then $H_{(n)}=H$, for some $n$. Then, $\widehat{H}^{(n)}=\mathbb{C}$ by Equation (8.4), and therefore ${\widehat{(H,\mathcal{B})}}$ is nilpotent. Idem for the converse with Equation (8.5).
@@ 8.2. Burnside property for nilpotent ARN-hypergroups
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup. Recall that we call $x_{i}\in\mathcal{B}$ a vanishing element if there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
@@
Let ${(H,\mathcal{B})}$ be an abelian RN-hypergroup and ${(L,\;\mathcal{S})}$ be a sub-hypergroup of ${(H,\mathcal{B})}$. Consider a standard element $x_{i}\in{\mathcal{B}}$ such that $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$. Then, $x_{i}$ is also vanishing element in ${(H,\mathcal{B})}$.
@@
Recall that $H//{\mathcal{S}}\simeq\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ as hypergroups, with $[x]\mapsto\lambda_{\mathcal{S}}x\lambda_{\mathcal{S}}$. Moreover, the hypergroup $\widehat{H//{\mathcal{S}}}$ can be identified with a subring of $\widehat{H}$. If $\lambda_{\mathcal{S}}=\sum_{j\in\mathcal{J}_{\mathcal{S}}}F_{j}$, then $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}=\oplus_{j\in\mathcal{J}_{\mathcal{S}}}\mathbb{C}F_{j}$, as a subalgebra of $H$. Thus, the characters $\mu_{j}$, with $j\in\mathcal{J}_{\mathcal{S}}$, have distinct restriction to $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, and cover all the characters of $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$. So if $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$, then $\lambda_{\mathcal{S}}x_{i}\lambda_{\mathcal{S}}$ is a vanishing element in $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, which in turn implies that $x_{i}$ is a vanishing element in $H$. ∎
@@
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. A grouplike element $g\in{G{(H,\mathcal{B})}}$ is a constituent of $x_{i}x_{i^{*}}$ if and only if $gx_{i}=\mathrm{FPdim}(g)x_{i}$. The set of all grouplike elements that are constituents of $x_{i}x_{i^{*}}$ form a subgroup of ${G{(H,\mathcal{B})}}$.
@@
Equation (5.1) gives that
$m(g,x_{i}x_{i}^{*})=\tau(x_{i}x_{i^{*}}g^{*})=\tau(x_{i^{*}}g^{*}x_{i})=m(x_{i^{*}},x_{i^{*}}g^{*})=m(x_{i},gx_{i}).$
Thus, by Lemma 2.27 one has $g^{-1}x_{i}=\mathrm{FPdim}(g^{-1})x_{i}$ if $m(g,x_{i}x_{i}^{*})>0$. The second assertion is a straightforward consequence of this. ∎
Proof of Theorem 1.12 It is enough to prove the following:
@@
If ${(H,\mathcal{B})}$ is of nilpotency class $1$, then ${(H,\mathcal{B})}$ is pointed and we are done.
@@
Consider $[x_{i}]\in H//G(H)$. Since $H//G(H)\simeq{(\widehat{H})_{ad}}$, it is nilpotent, with nilpotency class strictly less that of $H$, by Theorem 1.11. By induction hypothesis, we know that $[x_{i}]$ is either a grouplike element or a vanishing element in $H//G(H)$.
@@
Now, if $[x_{i}]$ is a grouplike element of $H//G(H)$, then
$\lambda_{G(H)}h_{[x_{i}]}x_{i}x_{i}^{*}\lambda_{G(H)}=\lambda_{G(H)},$
which implies that $h_{[x_{i}]}x_{i}x_{i}^{*}\in G(H)$. Thus, $h_{[x_{i}]}x_{i}x_{i}^{*}=\sum_{g\in G_{1}}g$, for a subgroup $G_{1}\subset G$. Since $x_{i}$ is not a grouplike element in $H$, $G_{1}\neq\{1\}$. Then, there is $\mu_{j}$ such that $\mu_{j}(\Lambda_{G_{1}})=0$ where $\Lambda_{G_{1}}=\frac{1}{|G_{1}|}\big{(}\sum_{x\in G_{1}}x\big{)}$ is the integral of $G_{1}$. This implies that $\mu_{j}(x_{i}x_{i}^{*})=0$. Thus $\mu_{j}(x_{i})=0$, which means that $x_{i}$ is a vanishing element. ∎

### v11

@@ 7.4. Computation of the regular components
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. For all $\mu_{j}\in{G(\widehat{H})}$,
(7.4) $F_{j}=\frac{1}{n(H)}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}.$
@@
For all $\mu_{j}\in{G(\widehat{H})}$, by Equation (2.4):
$F_{j}$ $=$ $\frac{1}{n_{j}}\sum_{i\in\mathcal{I}}h_{i}\mu_{j}(x_{i^{*}})x_{i}=\frac{1}{n_{j}}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\sum_{x_{i}\in\mathcal{B}_{\psi}}h_{i}\mu_{j}(x_{i^{*}})x_{i}=$
$=$ $\frac{1}{n_{j}}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi^{-1}(\mu_{j})R_{\psi}=\frac{1}{n_{j}}\sum_{\psi\in\widehat{{G(\widehat{H})}}}\psi(\mu_{j})R_{\psi^{-1}}.$
Finally, if $\mu_{j}\in{G(\widehat{H})}$ then $n_{j}=n(H)$ by Corollary 2.29. ∎
@@
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. The set $\mathcal{B}_{\psi}$ is not empty.
@@
Lemma 7.12 shows that the linear span of $\{F_{j}\;|\;\mu_{j}\in{G(\widehat{H})}\}$ lies inside the span of $\{\mathcal{R}_{\psi}\;|\;\psi\in\widehat{{G(\widehat{H})}}\text{ with }\mathcal{B}_{\psi}\neq\emptyset\}$. But $\widehat{{G(\widehat{H})}}\simeq{G(\widehat{H})}$ as finite abelian group, so have same order. Thus, an argument involving the dimension of these vector spaces implies that $\mathcal{B}_{\psi}$ is not empty, for any $\psi\in\widehat{{G(\widehat{H})}}$. ∎
By Equation (7.4) and the second orthogonality relation (2.6) applied on the (hyper)group ${G(\widehat{H})}$, it follows that:
@@
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Then
$R_{\psi}=\frac{n(H)}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}\psi(\mu_{j})F_{j}\big{)}.$
Proof of Theorem 7.8 and Theorem 7.9
By Lemma 7.14, and Corollary 6.6, if $\psi=1$, then
$R_{1}=\frac{n(H)}{|{G(\widehat{H})}|}\big{(}\sum_{\mu_{j}\in{G(\widehat{H})}}F_{j}\big{)}=\frac{n(H)}{|{G(\widehat{H})}|}\lambda_{{H_{ad}}}.$
Proposition 6.12 proves that $\mathcal{B}_{1}=\mathcal{B}_{ad}$. Therefore the above grading
@@
coincides with the universal grading of ${(H,\mathcal{B})}$. Thus $U_{H}={G(\widehat{H})}$. By duality, $U_{\widehat{H}}=G(H)$ if $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
7.5. On the perp of the adjoint sub-hypergroup
Recall from Definition 1.6 that an ARN-hypergroup is called dualizable if $^{\widehat{(H,\mathcal{B})}}$ is also (A)RN.
For any sub-hypergroup ${(L,\mathcal{S})}$ of a dualizable ARN-hypergroup ${(H,\mathcal{B})}$, define
@@ 7.5. On the perp of the adjoint sub-hypergroup
With the above notations, $\mathcal{S}^{\perp}=\bigcap_{s\in\mathcal{S}}\ker_{(H,\mathcal{B})}(s)$. Moreover, [Har79, Proposition 2.11] implies that $\mathcal{S}^{\perp}$ is a sub-hypergroup of $^{\widehat{(H,\mathcal{B})}}$, and $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$.
@@
Let $H$ be a dualizable ARN-hypergroup. With the above notations,
@@
Thus, by $(\mathcal{S}^{\perp})^{\perp}=\mathcal{S}$,
${G(\widehat{H})}=(H_{ad})^{\perp}\text{ and }G(H)={(\widehat{H})_{ad}}^{\perp}.$
@@
Consider the trivial character $\psi_{1}=x_{1}\big{|}_{G(\widehat{H})}$. It follows that
@@
By duality, we obtain the second equality. ∎
@@ 7.6. Quotients of hypergroups
In this subsection, we define the quotient hypergroup of a RN-hypergroup ${(H,\mathcal{B})}$ by a sub-hypergroup ${(L,\mathcal{S})}$. This is the analogue of the quotient construction from [Har79], in the case of probability groups. Recall that a probability group is a normalized RN-hypergroup.
We can define the following equivalence relation on $\mathcal{B}$. For two elements $a,b\in\mathcal{B}$, say $a\sim_{\mathcal{S}}b$ if and only if there are $s_{1},s_{2}\in{\mathcal{S}},x\in{\mathcal{B}}$ such that $m(x,as_{1})>0$ and $m(x,s_{2}b)>0$.
Let $[a]_{\mathcal{S}}$ denote the equivalence class of any $a\in\mathcal{B}$, with respect to $\sim_{\mathcal{S}}$. Define $H//{\mathcal{S}}$ as the set of all these equivalence classes of $\sim_{\mathcal{S}}$. Recall that $\lambda_{\mathcal{S}}$ is the central idempotent of ${\mathcal{S}}$ corresponding to $\mu_{1}=\mathrm{FPdim}$.
For two elements $a,b\in\mathcal{B}$, we can see that $[a]_{\mathcal{S}}=[b]_{\mathcal{S}}$ if and only if $\lambda_{\mathcal{S}}a\lambda_{\mathcal{S}}=\lambda_{\mathcal{S}}b\lambda_{\mathcal{S}}$. Therefore there is a set bijection
@@
Then, it is easy to verify that $H//{\mathcal{S}}$ becomes a RN-hypergroup, with the multiplication inherited from $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, via the above isomorphism. We denote by $\overline{m}({[c]},\;[a][b])$ the multiplicity structure of $H//{\mathcal{S}}$. Therefore,
@@
We write shortly $[a]$ instead of $[a]_{\mathcal{S}}$ when no confusion is possible. Following [Har79], if ${(H,\mathcal{B})}$ is abelian, we can show that
@@ 7.7. Applying Harrison’s results
In this subsection, we also use the short notation $H$ (or $\mathcal{B}$) for a hypergroup $(H,\mathcal{B})$, if $\mathcal{B}$ (or $H$) are implicitly understood. For any sub-hypergroup ${(L,\mathcal{S})}$ of a dualizable ANR-hypergroup ${(H,\mathcal{B})}$, note that [Har79, Proposition 2.11] implies that there is an isomorphism of hypergroups:
@@
Dually, for ${\mathcal{S}}={(\widehat{H})_{ad}}$, we obtain that
@@ 8. On lower and upper central series of hypergroups
Let ${(L,\mathcal{S})}$ be a sub-hypergroup of an ARN-hypergroup ${(H,\mathcal{B})}$. Let ${\mathcal{S}}^{co}$ be the set of standard elements $x\in\mathcal{B}$ such that $xx^{*}\in\mathcal{{\mathcal{S}}}$. The commutator ${(L,\mathcal{S})}^{{co}}$ of ${(L,\mathcal{S})}$ in ${(H,\mathcal{B})}$ as the sub-hypergroup generated by ${\mathcal{S}}^{co}$.
@@
for any sub-hypergroup $\mathcal{S}$ of $\mathcal{B}$. By applying $\perp$ to (8.1), we obtain
@@
For an ARN-hypergroup ${(H,\mathcal{B})}$, define $H_{(0)}=\mathbb{C}$ and $H_{(n)}=(H_{(n-1)})^{{co}}$, for all $n\geq 1$. Then,
$H_{(1)}=\mathbb{C}^{{co}}=G{(H,\mathcal{B})}=H_{pt}.$
@@
Similarly to [GN08, Theorem 4.16], we can prove that $H^{(n)}=\mathbb{C}$ if and only if $H_{(n)}=H$, where $H$ is an ARN-hypergroup.
@@
An ARN-hypergroup is termed nilpotent if there exists some $n\in\mathbb{N}$ such that $H^{(n)}=\mathbb{C}$. The smallest such $n$ for which this condition is satisfied is referred to as the nilpotency class of $H$.
@@
It reduces to show that $\mu_{j}\star\mu_{j^{\#}}\in({\mathcal{S}}^{co})^{\perp}$, for all $\mu_{j}\in{\mathcal{S}}^{\perp}$.
@@
(8.4) $\widehat{H}^{(n)}\subseteq(H_{(n)})^{\perp}$
@@
$\widehat{H}^{(0)}=\widehat{H}=\mathbb{C}^{\perp}=H_{(0)}^{\perp}.$
@@
By duality, we obtain the second inclusion. ∎
@@
If ${(H,\mathcal{B})}$ is nilpotent then $H_{(n)}=H$, for some $n$. Then, $\widehat{H}^{(n)}=\mathbb{C}$ by Equation (8.4), and therefore $^{\widehat{(H,\mathcal{B})}}$ is nilpotent. Idem for the converse with Equation (8.5).
@@ 8.2. Burnside property for nilpotent ARN-hypergroups
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. Recall that we call $x_{i}\in\mathcal{B}$ a vanishing element if there is $\mu_{j}\in{\widehat{\mathcal{B}}}$ such that $\mu_{j}(x_{i})=0$.
@@
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup and ${(L,\mathcal{S})}$ be a sub-hypergroup of ${(H,\mathcal{B})}$. Consider a standard element $x_{i}\in{\mathcal{B}}$ such that $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$. Then, $x_{i}$ is also a vanishing element in ${(H,\mathcal{B})}$.
@@
Recall that $H//{\mathcal{S}}\simeq\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ as hypergroups, with $[x]\mapsto\lambda_{\mathcal{S}}x\lambda_{\mathcal{S}}$. If $\lambda_{\mathcal{S}}=\sum_{j\in\mathcal{I}_{\mathcal{S}}}F_{j}$ (Definition 6.3), then $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}=\oplus_{j\in\mathcal{I}_{\mathcal{S}}}\mathbb{C}F_{j}$, as a subalgebra of $H$. Thus, the characters $\mu_{j}$, with $j\in\mathcal{I}_{\mathcal{S}}$, have distinct restriction to $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ (since $\mu_{i}(F_{j})=\delta_{i,j}$), and so cover all the characters of $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$ (since $\dim_{\mathbb{C}}(\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}})=|I_{\mathcal{S}}|$). If $[x_{i}]$ is a vanishing element in $H//{\mathcal{S}}$, then $\lambda_{\mathcal{S}}x_{i}\lambda_{\mathcal{S}}$ is a vanishing element in $\lambda_{\mathcal{S}}H\lambda_{\mathcal{S}}$, thus by above, $x_{i}$ is a vanishing element in $H$. ∎
@@
Let ${(H,\mathcal{B})}$ be an ARN-hypergroup. A grouplike element $x_{i}\in{G{(H,\mathcal{B})}}$ is a constituent of $x_{j}x_{j^{*}}$ if and only if $x_{i}x_{j}=\mathrm{FPdim}(x_{i})x_{j}$. The set of all grouplike elements that are constituents of $x_{j}x_{j^{*}}$ form a subgroup of ${G{(H,\mathcal{B})}}$.
@@
By Frobenius reciprocity (2.18), $N_{j,j^{*}}^{i}=\frac{h_{j}}{h_{i}}N_{i,j}^{j}$, thus $x_{i}$ is a constituent of $x_{j}x_{j^{*}}$ if and only if $x_{j}$ is a constitutent of $x_{i}x_{j}$, if and only if (by Lemma 2.34) $\frac{x_{i}}{d_{i}}\frac{x_{j}}{d_{j}}=\frac{x_{j}}{d_{j}}$, meaning that $x_{i}x_{j}=\mathrm{FPdim}(x_{i})x_{j}$. The second assertion is a straightforward consequence of this. ∎
Proof of Theorem 1.15 By Theorem 1.14, it is enough to prove the following:
@@
If ${(H,\mathcal{B})}$ is of nilpotency class 111, then ${(H,\mathcal{B})}$ is pointed and we are done.
@@
Consider $[x_{i}]\in H//G(H)$. Since $H//G(H)\simeq\widehat{{(\widehat{H})_{ad}}}$ by (7.7), it is nilpotent, with nilpotency class $n-1$, by Theorem 1.14. By induction hypothesis, we know that $[x_{i}]$ is either a grouplike element or a vanishing element in $H//G(H)$.
@@
Now, if $[x_{i}]$ is a grouplike element of $H//G(H)$ and by abelian assumption,
$\lambda_{G(H)}=h_{[x_{i}]}\lambda_{G(H)}x_{i}\lambda_{G(H)}x_{i}^{*}\lambda_{G(H)}=\lambda_{G(H)}h_{[x_{i}]}x_{i}x_{i}^{*}\lambda_{G(H)},$
which implies that $h_{[x_{i}]}x_{i}x_{i}^{*}\in G(H)$. Thus, $h_{[x_{i}]}x_{i}x_{i}^{*}=\sum_{g\in G_{1}}g$, for a subgroup $G_{1}\subset G$ (by Lemma 8.8). Since $x_{i}$ is not a grouplike element in $H$, $G_{1}\neq\{1\}$. Then, there is $\mu_{j}$ such that $\mu_{j}(\Lambda_{G_{1}})=0$ where $\Lambda_{G_{1}}=\frac{1}{|G_{1}|}\big{(}\sum_{x\in G_{1}}x\big{)}$ is the integral of $G_{1}$. This implies that $\mu_{j}(x_{i}x_{i}^{*})=0$. Thus $\mu_{j}(x_{i})=0$, which means that $x_{i}$ is a vanishing element. ∎